The residents of an agrarian country spend all their income on cauliflower, cucumbers, and carrots. In 2017, they produced 20 heads of cauliflower valued at Rs.200, 5 kgs of cucumbers for Rs.75, and 50 carrots for Rs.80. In 2018, they produced 25 heads of cauliflower for Rs.250, 8 kgs of cucumbers valued at Rs.160, and 50 carrots valued at Rs.150. a. Calculate nominal and real GDPs taking 2018 as the base year. Then calculate GDP-deflator index. b. Assuming that a common person consumes 1 heads of cauliflower, two kg. of cucumbers and 5 carrots, calculate CPI taking 2018 as the base year. c. Calculate inflation rate between 2017-18 measured both from GDP-deflator index and CPI?


1
Expert's answer
2021-04-26T19:35:23-0400

a. Nominal and real GDPs taking 2018 as the base year are:

NGDP(2017) = 200×20 + 75×5 + 80×50 = 8,375,

RGDP(2017) = 250×20 + 160×5 + 150×50 = 13,300,

NGDP(2018) = 250×25 + 160×8 + 150×50 = 15,030,

RGDP(2018) = 15,030.

GDP-deflator index = 8,375/13,300 = 0.63.

b. CPI(2017) = (200×1 + 75×2+ 80×5)/(250×1 + 160×2 + 150×5)×100 = 750/1320×100 = 56.8.

c. Inflation rate between 2017-18 measured from GDP-deflator index is:

0.63 - 1 = -0.37 or -37%,

and from CPI is:

56.8 - 100 = -43.2%.